OpenStudy (anonymous):

What is the largest common denominator of 4/9 and 7/12?

OpenStudy (anonymous):

108 I believe

OpenStudy (amistre64):

nope

OpenStudy (amistre64):

its a common one, but not the most efficient

OpenStudy (anonymous):

sorry bad lag :/

OpenStudy (anonymous):

ok sorry

OpenStudy (amistre64):

a good way to find the LCD is to math it out: \[\frac19+\frac1{12}=n\] and start with the largest denominator and multiply it thru \[\frac{\cancel{12}^4}{\cancel{9}^3}+\frac{12}{12}=12n\] \[\frac43+1=12n\]and do it again \[\frac{12}3+1(3)=12(3)n\] \[4+1(3)=\underbrace{36}_{LCD}~n\]

OpenStudy (anonymous):

ty that's what I thought it was!

OpenStudy (amistre64):

spose to be lowest common denom .... largest? whew ... it could be in the billions of billions